# Unknotting Lagrangian $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $\mathbb{R}^{2n}$

### v1

@@ 1. Introduction and overview
The purpose of this paper is to describe the isotopy classes represented by Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$. The case $n=1$ is elementary, so we assume $n\geq 2$. $\mathbb{R}^{2n}$ is identified with $\mathbb{C}^{n}$ so that the symplectic form is given by the imaginary part of the standard Hermitian metric. An embedding $j:N\to\mathbb{C}^{n}$ is Lagrangian if multiplication by $i=\sqrt{-1}$ maps the tangent space of $j(N)$ at each point onto its Euclidean orthogonal complement. Lagrangian embeddings are obviously totally real, that is, have no tangent complex lines.
@@
$\mathrm{S}^{1}\times\mathrm{S}^{n-1}\ni(\theta,x)\longmapsto(1+e^{i\theta}/2)\bigl{(}x_{1},...,x_{n}\bigr{)}\in\mathbb{C}^{n},$ (1.1)
@@
Our main Theorem 1.1 shows that for $n\neq 2,4$ every Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is isotopic to (1.1) if $n$ is even and either to (1.1) or to its composition with the reflection $(\theta,x)\mapsto(-\theta,x)$ if $n$ is odd. The isotopy is not through Lagrangian embeddings in general. In fact, there exist Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ for all $n>2$ that are not isotopic through Lagrangian embeddings to any reparametrisation of (1.1), see [18, Theorem 5] and [8, Corollary 1.6].
@@
The result for $n=8$ was obtained by Borrelli [3]; our argument is a generalisation of his approach based on [17, §6]. Dimitroglou Rizell and Evans [5, Corollary E] proved the result for Lagrangian embeddings with minimal Maslov number $n$ for even $n\geq 6$ and for monotone Lagrangian embeddings for odd $n\geq 5$. However, embeddings (1.1) have minimal Maslov number $2$ for all $n$ and non-monotone Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{2k}\to\mathbb{C}^{2k+1}$ exist for all $k\geq 1$ by [8, Corollary 1.6].
@@
All Lagrangian submanifolds in $\mathbb{C}^{n}$ diffeomorphic to $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ are smoothly isotopic if $n\neq 4$.
@@
Theorem 1.1 is a typical ‘symplectic rigidity’ result. It becomes completely false if Lagrangian embeddings are replaced by ‘soft’ totally real embeddings in $\mathbb{C}^{n}$. Namely, every embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is smoothly isotopic to a totally real embedding. To see this, recall that any two embeddings of an orientable $n$-manifold in $\mathbb{C}^{n}$ are regularly homotopic through immersions. (This follows from the general theory in [14].) Therefore the tangent map of any embedding of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ is homotopic to the totally real tangent map of an (existing) Lagrangian embedding. An isotopy to a totally real embedding is then given by the $h$-principle [11, §19.3].
@@
The proof of Theorem 1.1 follows the scheme from [3] and consists of two parts discussed in Sections 2 and 3 respectively:
@@
(1) For a totally real embedding $j:\mathrm{S}^{1}\times\mathrm{S}^{n-1}\to\mathbb{C}^{n}$ and a nowhere vanishing vector field $v$ tangent to the $\mathrm{S}^{1}$-fibres, we may consider the pushoff of $j$ in the direction of $i\cdot dj(v)$. If the embedding is Lagrangian, this pushoff is nullhomologous in $\mathbb{C}^{n}\setminus j(\mathrm{S}^{1}\times\mathrm{S}^{n-1})$ or, equivalently, is unlinked with $(n-1)$-cycles in $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ up to replacing $v$ with $-v$ for odd $n$. The proof of this assertion uses a generalisation of the Luttinger surgery [15] to arbitrary dimensions in the same way as the original surgery was used in [10]. Such a surgery produces (for $n\geq 3$) an exact symplectic manifold that is symplectomorphic to $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$ outside of a compact subset and must therefore have the homology of $\mathbb{R}^{2n}$ by a basic application of pseudoholomorphic curves [16]. Comparing this with the homology calculations in the explicit model from [17, §6] proves the vanishing of the relevant linking numbers.
@@
(2) Embeddings of an $n$-manifold $N$ in $\mathbb{R}^{2n}$ for $n\geq 4$ are governed by an $h$-principle established by Haefliger and Hirsch [13]. Isotopy classes of such embeddings are in one-to-one correspondence with homotopy classes of certain equivariant maps from the frame bundle of $N^{\circ}$ to the Stiefel manifold $\mathcal{V}_{2n,n+1}$, where $N^{\circ}$ is the complement to a small ball in $N$. The additional $(n+1)$-st component of the map is given by a normal vector field such that the pushoff of $N$ is unlinked with $N^{\circ}$. In particular, for a Lagrangian embedding of $N=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$, this normal field can be taken to be $i\cdot dj(\pm v)$ by part (1). (The sign only makes a difference for odd $n$ and can be adjusted by a reflection of $\mathrm{S}^{1}$.) If we now recall that our manifold is parallelisable and our embedding is totally real, the relevant frame map reduces to a map from $N^{\circ}$ itself of the form $x\mapsto\Phi(x)\gamma(x)$, where $\Phi(x)\in\mathrm{GL}(n,\mathbb{C})$ and $\gamma$ is a fixed map to the Stiefel manifold. Homotopy calculations (conveniently done in [5]) show that for $n\neq 2,4$ the map $\Phi$ has no effect on the homotopy class, and the result for $n>4$ follows. In the remaining case $n=3$, a version of the same argument in the setting of [19] determines the value of the Whitney invariant $W$, which turns out to be ±$\pm 1$, so that there is only one possible isotopy class up to a reflection of $\mathrm{S}^{1}$.
@@ 2. Generalised Luttinger surgery, symplectic rigidity, and the linking class
Take the orientation on $\mathbb{R}^{n}$ by the order of the coordinates and orient $T^{*}\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ as $\mathbb{R}^{n}\oplus i\mathbb{R}^{n}$. (This orientation differs from the standard complex/symplectic one by the sign $(-1)^{\frac{1}{2}n(n-1)}$, which will make some of the following formulas nicer.)
@@
Let $f$ and $\tau$ be the diffeomophisms of $S_{r}^{*}\Sigma_{n}$ defined above. Then
@@
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be a Lagrangian embedding and take the nowhere vanishing vector field $v=\sum x_{k}\frac{\partial}{\partial x_{k}}$ on $\Sigma_{n}=\mathbb{R}_{x}^{n}\setminus\{0\}/\mathbb{Z}$.
@@
Note also that by (2.6), (2.5), and the (anti-)commutation rule for the linking number (see e.g. [7, p. 234]), we have
@@ 3. Classification of embeddings
Let us start with the special case $n=3$. The isotopy classes of smooth embeddings $j:\Sigma_{3}\to\mathbb{R}^{6}\cong\mathbb{C}^{3}$ are indexed by the following two invariants [19, p. 648]. First, the Whitney invariant
@@
$2W(j)=\pm\mathop{\mathrm{PD}}\,e(\xi^{\perp}))\in\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z}),$ (3.1)
@@
For a Lagrangian embedding $j$, an unlinked normal bundle section is given by $\xi:=i\cdot dj(v)$ after possibly precomposing $j$ with the diffeomorphism (2.4) by Corollary 2.9. Multiplication by $i$ defines an isomorphism $\xi^{\perp}\cong v^{\perp}$, where the latter complement is taken in $T\Sigma_{3}$. The orthogonal complement to $v$ is just the subbundle tangent to the $\mathrm{S}^{2}$-fibres in $\Sigma_{3}$, so its Euler class is clearly equal to twice a generator of $\mathrm{H}^{2}(\Sigma_{3};\mathbb{Z})$, whence
@@
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be an embedding and let $\Sigma_{n}^{\circ}$ denote the complement of a small $n$-ball in $\Sigma_{n}$. There exists a unique up to homotopy normal vector field $\nu_{j}$ on $j(\Sigma_{n})$ such that it does not vanish on $\Sigma_{n}^{\circ}$ and the pushoff of $j(\Sigma)$ by $\nu_{j}$ is nullhomologous in $\mathbb{C}^{n}\setminus j(\Sigma_{n}^{\circ})$, see [13, p. 134]. (This vector field is called the associated vector field of $j$ in [3, §5.1] and the Haefliger–Hirsch vector field in [5, §6.2]. It is also the unlinked section of the normal bundle from [19] appearing in (3.1).)
@@
to the Stiefel manifold of real $(n+1)$-frames in $\mathbb{R}^{2n}\cong\mathbb{C}^{n}$, cf. [3, §5.2]. (It is convenient to use arbitrary frames here but it is well-known that $\mathcal{V}_{m,\ell}$ retracts onto the Stiefel manifold of orthonormal frames [20, §2].)
Since $\mathcal{V}_{2n,n+1}$ is $(n-2)$-connected [20, Satz 8] and $\Sigma_{n}^{\circ}$ retracts onto the wedge sum $\mathrm{S}^{1}\vee\mathrm{S}^{n-1}$, it follows that for $n\geq 3$ the homotopy class of (3.2) is determined by the image of the meridional $(n-1)$-sphere $m\subset\Sigma_{n}^{\circ}$ in $\pi_{n-1}(\mathcal{V}_{2n,n+1})$. The latter group is isomorphic to $\mathbb{Z}$ if $n$ is odd and to $\mathbb{Z}/2\mathbb{Z}$ if $n$ is even [20, Satz 9], which completes the classification of smooth embeddings of $\Sigma_{n}$ in $\mathbb{C}^{n}$ for $n\geq 4$.
@@
$x\longmapsto\left(dj_{x}(\zeta_{1}(x)),\dots,dj_{x}(\zeta_{n}(x)),i\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}dj_{x}(\zeta_{k}(x))\right)$
@@
where $\{e_{k}\}_{k=1,\dots,n}$ is our fixed complex frame of $\mathbb{C}^{n}$. Hence, the map (3.2) can be written as
$x\longmapsto\left(\Phi(x)e_{1},\dots,\Phi(x)e_{n},\Phi(x)\left(\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right)\right).$ (3.3)
The action of $\mathrm{GL}(n,\mathbb{C})\subset\mathrm{GL}_{+}(2n,\mathbb{R})$ on $\mathcal{V}_{2n,n+1}$ induces a trivial map on the relevant homotopy group $\pi_{n-1}$ for all $n\neq 2,4$ by the proof of [5, Lemma 6.14]. Specifically,
@@
If $n=4$, the following argument from [3, p. 232] shows that both isotopy classes of embeddings $j:\Sigma_{4}\to\mathbb{C}^{4}$ have Lagrangian representatives. Note first that (for any $n\geq 3$) the action of the map $\Phi$ in (3.3) on $\pi_{n-1}(\Sigma_{n}^{\circ})$ can be chosen at will by [1, §6.2] and the $h$-principle for Lagrangian immersions of $\mathrm{S}^{n-1}$. However, the map
@@
with the $+$ sign for $j$ and the -- sign for $j\circ\rho$. The restrictions of the maps (3.5) to the meridional sphere $m=\{\|x\|=1\}$ parametrise the (homotopy) fibre $\mathrm{S}^{n-1}\sim\mathcal{V}_{n,1}$ of the ‘forgetful’ fibre bundle
@@
and differ by the antipodal map of $\mathrm{S}^{n-1}$. For odd $n$, the antipodal map is orientation reversing and hence acts non-trivially on $\pi_{n-1}(\mathrm{S}^{n-1})=\mathbb{Z}$. Furthermore, $\pi_{n-1}(\mathcal{V}_{2n,n})=0$ by [20, Satz 8] and $\pi_{n}(\mathcal{V}_{2n,n})=\mathbb{Z}/2\mathbb{Z}$ for odd $n>1$ by [20, Satz 9], so the homotopy exact sequence
@@
Remark 3.5.
For $n=3$ one can alternatively (and in fact equivalently) use formula (3.1) and the specific choice of the unlinked section $\xi$ in the proof of Proposition 3.1 to show that $W(j\circ\rho)=-W(j)$ for a Lagrangian embedding $j$.
Theorem 1.1 follows by combining Propositions 3.1, 3.2, and 3.4.

### v3

@@ 1. Introduction and overview
The purpose of this paper is to describe the isotopy classes represented by Lagrangian embeddings of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ in $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$. We assume $n\geq 2$ to exclude the elementary special case $n=1$. $\mathbb{R}^{2n}$ is identified with $\mathbb{C}^{n}$ so that the symplectic form is the imaginary part of the standard Hermitian metric. An embedding $j:N\to\mathbb{C}^{n}$ is Lagrangian if multiplication by $i=\sqrt{-1}$ maps the tangent space of $j(N)$ at each point onto its Euclidean orthogonal complement. Lagrangian embeddings are obviously totally real, that is, have no tangent complex lines.
@@
$\mathrm{S}^{1}\times\mathrm{S}^{n-1}\ni\;(\theta,x_{1},\dots,x_{n})\longmapsto\bigl{(}1+\frac{1}{2}e^{i\theta}\bigr{)}(x_{1},\dots,x_{n})\;\in\mathbb{C}^{n},$ (1.1)
@@
Our main Theorem 1.1 shows that for $n\neq 2,4$ every Lagrangian embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is isotopic to (1.1) if $n$ is even and either to (1.1) or to its composition with the reflection $(\theta,x)\mapsto(-\theta,x)$ if $n$ is odd. The isotopy is not through Lagrangian embeddings in general. In fact, there exist Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ for all $n>2$ that are not isotopic through Lagrangian embeddings to any reparametrisation of (1.1), see [18, §4], [9, Corollary 1.6], and [19, §4].
@@
The result for $n=8$ was obtained by Borrelli [3]; our argument is a generalisation of his approach based on [17, §6]. Dimitroglou Rizell and Evans [6, Corollary E] proved the result for Lagrangian embeddings with minimal Maslov number $n$ for even $n\geq 6$ and for monotone Lagrangian embeddings for odd $n\geq 5$. However, embeddings (1.1) have minimal Maslov number $2$ for all $n$ and non-monotone Lagrangian embeddings $\mathrm{S}^{1}\times\mathrm{S}^{2k}\to\mathbb{C}^{2k+1}$ exist for all $k\geq 1$ by [9, Corollary 1.6].
@@
If $n\neq 4$, then all Lagrangian submanifolds in $\mathbb{C}^{n}$ diffeomorphic to $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ are smoothly isotopic to the image of (1.1).
@@
Theorem 1.1 is a typical ‘symplectic rigidity’ result. It becomes completely false if Lagrangian embeddings are replaced by ‘soft’ totally real embeddings in $\mathbb{C}^{n}$. Namely, every embedding $\mathrm{S}^{1}\times\mathrm{S}^{n-1}\hookrightarrow\mathbb{C}^{n}$ is smoothly isotopic to a totally real embedding. To see this, recall that any two embeddings of an orientable $n$-manifold in $\mathbb{C}^{n}$ are regularly homotopic through immersions. (This follows from the general theory in [14].) Therefore the tangent map of any embedding of $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ is homotopic to the totally real tangent map of an (existing) Lagrangian embedding. An isotopy to a totally real embedding is then given by the $h$-principle [4, §27.4].
@@
The proof of Theorem 1.1 follows the scheme from [3] and consists of two parts discussed in Section 2 and Section 3, respectively:
@@
(1) For a totally real embedding $j:\mathrm{S}^{1}\times\mathrm{S}^{n-1}\to\mathbb{C}^{n}$ and a nowhere vanishing vector field $v$ tangent to the $\mathrm{S}^{1}$-fibres, we may consider the pushoff of $j$ in the direction of $i\cdot dj(v)$. If the embedding is Lagrangian, this pushoff is nullhomologous in $\mathbb{C}^{n}\setminus j(\mathrm{S}^{1}\times\mathrm{S}^{n-1})$ or, equivalently, is unlinked with $(n-1)$-cycles in $\mathrm{S}^{1}\times\mathrm{S}^{n-1}$ up to replacing $v$ with $-v$ for odd $n$. The proof of this assertion uses a generalisation of the Luttinger surgery [15] to arbitrary dimensions in the same way as the original surgery was used in [11]. Such a surgery produces (for $n\geq 3$) an exact symplectic manifold that is symplectomorphic to $(\mathbb{R}^{2n},\omega_{\mathrm{st}})$ outside of a compact subset and that must therefore have the homology of $\mathbb{R}^{2n}$ by a basic application of pseudoholomorphic curves [16]. Comparing this with the homology calculations in the explicit model from [17, §6] proves the vanishing of the relevant linking numbers.
@@
(2) Embeddings of a connected orientable $n$-manifold $N$ in $\mathbb{R}^{2n}$ for $n\geq 4$ are governed by an $h$-principle due to Haefliger and Hirsch [13]. Isotopy classes of such embeddings are in one-to-one correspondence with homotopy classes of equivariant maps from the frame bundle of $N^{\circ}$ to the Stiefel manifold $\mathcal{V}_{2n,n+1}$, where $N^{\circ}$ is the complement to a small ball in $N$. The $(n+1)$-st component of the map to $\mathcal{V}_{2n,n+1}$ is given by a normal vector field on $N$ such that the pushoff of $N$ is unlinked with $N^{\circ}$. Thus, for a Lagrangian embedding of $N=\mathrm{S}^{1}\times\mathrm{S}^{n-1}$, this normal field can be taken to be $i\cdot dj(\pm v)$ by part (1). (The sign only makes a difference for odd $n$ and can be adjusted by a reflection of $\mathrm{S}^{1}$.) Recalling now that our manifold is parallelisable and our embedding is totally real, we can reduce the frame map to a map from $N^{\circ}$ itself of the form $x\mapsto\Phi(x)\beta(x)$, where $\Phi(x)\in\mathrm{GL}(n,\mathbb{C})$ and $\beta$ is a fixed map to the Stiefel manifold. Homotopy calculations (conveniently done in [6]) show that for $n\neq 2,4$ the map $\Phi$ has no effect on the homotopy class, and the result for $n>4$ follows. In the remaining case $n=3$, a version of the same argument in the setting of [20] shows that the Whitney invariant $W$ equals ±$\pm 1$ so that $\mathbb{Z}/|W|\mathbb{Z}=\{0\}$ and there is only one possible isotopy class up to a reflection of $\mathrm{S}^{1}$.
@@ 2. Generalised Luttinger surgery, symplectic rigidity, and the linking class
Take the orientation on $\mathbb{R}^{n}$ by the order of the coordinates and orient $T^{*}\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ as $\mathbb{R}^{n}\oplus i\mathbb{R}^{n}$. (This orientation differs from the standard complex/symplectic one by the sign $(-1)^{n(n-1)/2}$, which will make some of the following formulas nicer.)
@@
Let $f$ and $\tau$ be the diffeomorphisms of $S_{r}^{*}\Sigma_{n}$ defined above. Then
@@
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be a Lagrangian embedding and take the nowhere vanishing vector field $v=\sum x_{k}\frac{\partial}{\partial x_{k}}$ on $\Sigma_{n}=\left(\mathbb{R}_{x}^{n}\setminus\{0\}\right)/\mathbb{Z}$.
@@
Note also that by (2.6), (2.5), and the (anti-)commutation rule for the linking number (see e.g. [5, p. 234]), we have
@@ 3. Classification of embeddings
Let us start with the special case $n=3$. The isotopy classes of smooth embeddings $j:\Sigma_{3}\to\mathbb{R}^{6}\cong\mathbb{C}^{3}$ are indexed by the following two invariants [20, p. 648]. First, the Whitney invariant
@@
$2W(j)=\pm\mathop{\mathrm{PD}}\,e(\xi^{\perp})\in\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z}),$ (3.1)
@@
For a Lagrangian embedding $j$, an unlinked normal bundle section is given by $\xi:=i\cdot dj(v)$ after possibly precomposing $j$ with the diffeomorphism (2.4) by Corollary 2.9. Multiplication by $i$ defines an isomorphism $\xi^{\perp}\cong v^{\perp}$, where the latter complement is taken in $T\Sigma_{3}$. The orthogonal complement to $v$ is just the subbundle tangent to the $\mathrm{S}^{2}$-fibres in $\Sigma_{3}$, so its Euler class equals twice a generator of $\mathrm{H}^{2}(\Sigma_{3};\mathbb{Z})$. Hence,
@@
Let $j:\Sigma_{n}\to\mathbb{C}^{n}$ be an embedding and let $\Sigma_{n}^{\circ}$ denote the complement of a small $n$-ball in $\Sigma_{n}$. There exists a unique up to homotopy normal vector field $\nu_{j}$ on $j(\Sigma_{n})$ such that it does not vanish on $\Sigma_{n}^{\circ}$ and the pushoff of $j(\Sigma_{n})$ by $\nu_{j}$ is nullhomologous in $\mathbb{C}^{n}\setminus j(\Sigma_{n}^{\circ})$, see [13, p. 134]. (This vector field is called the associated vector field of $j$ in [3, §5.1] and the Haefliger–Hirsch vector field in [6, §6.2]. It is also the unlinked section of the normal bundle from [20] appearing in (3.1).)
@@
to the Stiefel manifold of real $(n+1)$-frames in $\mathbb{R}^{2n}\cong\mathbb{C}^{n}$, cf. [3, §5.2]. (It is convenient to use arbitrary frames here but it is well-known that $\mathcal{V}_{*,*}$ retract onto Stiefel manifolds of orthonormal frames [21, §2].)
Since $\mathcal{V}_{2n,n+1}$ is $(n-2)$-connected [21, Satz 8] and $\Sigma_{n}^{\circ}$ retracts onto the wedge sum $\mathrm{S}^{1}\vee\mathrm{S}^{n-1}$, it follows that for $n\geq 3$ the homotopy class of (3.2) is determined by the image of the meridional $(n-1)$-sphere $m\subset\Sigma_{n}^{\circ}$ in $\pi_{n-1}(\mathcal{V}_{2n,n+1})$. The latter group is isomorphic to $\mathbb{Z}$ if $n$ is odd and to $\mathbb{Z}/2\mathbb{Z}$ if $n$ is even [21, Satz 9], which completes the classification of smooth embeddings of $\Sigma_{n}$ in $\mathbb{C}^{n}$ for $n\geq 4$.
@@
$x\longmapsto\left(dj_{x}(\zeta_{1}(x)),\dots,dj_{x}(\zeta_{n}(x)),i\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}dj_{x}(\zeta_{k}(x))\right).$
@@
where $e_{k}$, $k=1,\ldots,n$, is our fixed complex frame of $\mathbb{C}^{n}$. Hence, the map (3.2) can be written as
$x\longmapsto\left(\Phi(x)e_{1},\dots,\Phi(x)e_{n},\Phi(x)\sum\limits_{k=1}^{n}\frac{x_{k}}{\|x\|}ie_{k}\right).$ (3.3)
The action of $\mathrm{GL}(n,\mathbb{C})\subset\mathrm{GL}_{+}(2n,\mathbb{R})$ on $\mathcal{V}_{2n,n+1}$ induces a trivial map on the relevant homotopy group $\pi_{n-1}$ for all $n\neq 2,4$ by the proof of [6, Lemma 6.14]. Specifically,
@@
If $n=4$, the following argument from [3, p. 232] shows that both isotopy classes of embeddings $j:\Sigma_{4}\to\mathbb{C}^{4}$ have Lagrangian representatives. Note first that (for any $n\geq 3$) the action of the map $\Phi$ in (3.3) on $\pi_{n-1}(\Sigma_{n}^{\circ})$ can be chosen at will by [1, §6.2] and the $h$-principle for Lagrangian immersions of $\mathrm{S}^{n-1}$ [4, §24.3]. However, the map
@@
Remark 3.4.
For $n=3$, the proof of Proposition 3.2 amounts to an alternative computation of the Whitney invariant. Let $\mathcal{N}\to\mathcal{V}_{2n,n+1}$ be the tautological normal bundle. (The fibre of $\mathcal{N}$ at a frame is the orthogonal complement to the span of the frame.) An orientation on $\mathbb{R}^{2n}$ defines an orientation on $\mathcal{N}$. Two observations are immediate:
  1. (1)
    The pull-back of $\mathcal{N}$ by the Haefliger–Hirsch map (3.2) is the orthogonal complement $\nu_{j}^{\perp}(=\xi^{\perp})$ in the normal bundle of $j$.
  2. (2)
    The pull-back of $\mathcal{N}$ to the meridional $2$-sphere $m=\{\|x\|=1\}$ by the standard map (3.4) is the tangent bundle of $m$.
For a Lagrangian embedding $j$ (precomposed with (2.4) if needed), the maps (3.2) and (3.4) are homotopic. Hence, $\langle e(\xi^{\perp}),[m]\rangle=2$ and therefore $W(j)=\pm 1\in\mathbb{Z}\cong\mathrm{H}_{1}(\Sigma_{3};\mathbb{Z})$ by (3.1).
@@
with the $+$ sign for $j$ and the -- sign for $j\circ\rho$. The restrictions of the maps (3.5) to the meridional sphere $m=\{\|x\|=1\}$ parametrise the (homotopy) fibre $\mathrm{S}^{n-1}\sim\mathcal{V}_{n,1}\times\mathbb{R}^{n}$ of the ‘forgetful’ fibre bundle
@@
and differ by the antipodal map of $\mathrm{S}^{n-1}$. For odd $n$, the antipodal map is orientation reversing and hence acts non-trivially on $\pi_{n-1}(\mathrm{S}^{n-1})=\mathbb{Z}$. Furthermore, $\pi_{n-1}(\mathcal{V}_{2n,n})=0$ by [21, Satz 8] and $\pi_{n}(\mathcal{V}_{2n,n})=\mathbb{Z}/2\mathbb{Z}$ for odd $n>1$ by [21, Satz 9], so the homotopy exact sequence
@@
Theorem 1.1 follows by combining Propositions 3.1, 3.2, and 3.5.
Acknowledgements
The author is grateful to Leonid Polterovich, Felix Schlenk, and Arkadiy Skopenkov for their comments on the first draft of this paper.